Given the length of the vector and angle, determine the components of the vector. Magnitude (length) of A = 5.0990 Angle vector A makes with respect to the x-axis = 168.6901 degrees

1. Ax =

2. Ay =
Answer #1

here,

magnitude of vector A , |A| = 5.099 units

theta = 168.6901 degree

the x-compoenent of vector , Ax = A * cos(theta)

Ax = 5.099 * cos(168.6901)

Ax = - 5 units

the y-component of vector , Ay = A * sin(theta)

Ay = 5.099 * sin(168.6901)

Ay = 1 unit